OpenStudy (anonymous):

Create your own unique literal equation and solve for one of the variables. Show your work. Then, using complete sentences, explain how you solved for the variable you chose. (2 points)

OpenStudy (anonymous):

PV=nRT Solved for T: T=PV/(nR) Guiding Principle: Everything you do to one side of the equation must be reproduced on the other side. To solve for T, I divide both sides of the equation by the quantity nR PV/(nR)=nRT/(nR) The equation becomes PV/(nR)=T because the quantity (nR)/(nR)=1 The End